OpenStudy (anonymous):

write an equation for each line 1.) through (4, -1) with m=-3 2.) parallel to y=4 with x-intercept (-3, 0)

OpenStudy (anonymous):

1) line equation y=mx+b, m=-3 y=-3*x+b -1=-3*4+b -1=-12+b b=11 y=-3x+11

OpenStudy (anonymous):

Well given that the slope is -3. You have -3x + B = Y. Plug in x = 4 and y = -1. -3(4) + B = -1. Solve for B. Done with the first one. Second one is quite similar, except you have the same slope.

OpenStudy (anonymous):

Or you can just use the point-slope formula.

OpenStudy (anonymous):

2), someone correct me if im wrong here, but to be parallel to y=4 means the slope is 0. thus, to touch (-3,0), you must have an equation y=0

OpenStudy (anonymous):

which is the x axis so the intercept data is irrelevant...